OpenStudy (anonymous):

kainuiiiiiiiiiii, I need you to explain to me what exactly you did with this integral haha.

OpenStudy (anonymous):

@Kainui

OpenStudy (kainui):

i 4get sry

OpenStudy (kainui):

Haha, no seriously, what's up?

OpenStudy (anonymous):

\[\huge \int\limits_{-\infty}^{\infty}e ^{-x ^{2}}dx\]

OpenStudy (kainui):

Ahh, so now just multiply that same integral by itself and take the square root of it. Since the integral doesn't depend on x, make the other integral you multiply it by just with y's instead.

OpenStudy (kainui):

That's just to make it more obviously different from it if that makes any sense at all lol

OpenStudy (anonymous):

I see, so would this method work even without the limits?

OpenStudy (kainui):

I dunno, I tried and couldn't really figure it out.

OpenStudy (anonymous):

Ah, yeah this integral has been haunting me for quite some time..

OpenStudy (kainui):

You could try the more general version though: \[\int\limits _\infty ^\infty ae^{-bx^2}dx\] Then you could differentiate the answer you get with respect to a or b depending on what comes out. Actually I'll try that out now. =P

OpenStudy (kainui):

well obviously that integral up there =0 but w/e

OpenStudy (kainui):

You know what I meant lol

OpenStudy (anonymous):

Mhm wonder if you would still get...\[\frac{ 1 }{ a^2 }(au-1)e ^{au}\]

OpenStudy (anonymous):

And yeah lol

OpenStudy (anonymous):

+C

OpenStudy (kainui):

idk where did you get that 1/a^2 thing?

OpenStudy (anonymous):

Back of my calc textbook lol, I'm looking at tables of integrals :P

OpenStudy (kainui):

\[\sqrt{\pi}*a^{-1/2}=\int\limits_{-\infty}^{\infty} e^{-ax^2}dx\] This is what I'm getting so check this out

OpenStudy (kainui):

\[\frac{1}{2}\sqrt{\pi}*a^{-3/2}=\int\limits_{-\infty}^{\infty} x^2e^{-ax^2}dx\]